OpenStudy (anonymous):

Help please! The cost of a parking ticket at IU is $35 for the first offense, but the cost triples for each additional offense. Write a formula for the cost C as a function of the number of tickets n.

OpenStudy (anonymous):

Let's work it out. First offence(n=1): P=35 Second offence(n=2): P=35*3 Third offence(n=3): P=(35*3)*3=35*3² Fourth offence(n=4): P=(35*3²)*3=35*3³ Now we can see that the power of three each is n-1, so we can a general function for the cost, C: \[C=35\times3^{n-1}\]

OpenStudy (mathstudent55):

Ticket number Amount Power form 1 $35 $35 2 $35 * 3 $35 * 3^1 3 $35 * 3 * 3 $35 * 3^2 4 $35 * 3 * 3 * 3 $35 * 3^3 (notice exponent is 1 less than ticket number) ... ... ... n $35 * 3^(n - 1)

OpenStudy (anonymous):

Both of you are legends! Thanks for the help, it is greatly appreciated.

OpenStudy (mathstudent55):

You're welcome.